Phone calls that come to the customer service number are independent. Average waiting time until the fifth phone call is 10 minutes.

a.Find the average waiting time between any two consecutive phone calls.

b. Find the variance of the waiting time until the fifth phone call to customer service

.c. Find the variance of the waiting time until the first phone call to customer service.

d. Find the 80th percentile of the waiting time until the first phone call.
Answer #1

a) average waiting time between any two consecutive phone calls =10/5 =2

b)  variance of the waiting time until the fifth phone call to customer service =\alpha\beta2 =5*22 =20

c)

variance of the waiting time until the first phone call to customer service =\beta2 =22 =4

d)

80th percentile =-\beta*ln(1-p) =-2*ln(1-0.80)=3.22 minutes